# Bayesian Group Regularization in Generalized Linear Models with a Continuous Spike-and-Slab Prior thanks: Keywords and phrases: generalized linear model, group sparsity, spike-and-slab, spike-and-slab group lasso, posterior contraction, rate of convergence, variable selection

### v7

@@ 6.4 Fully Bayesian inference
  • Experiment 9 (grouped logistic regression). With $n=300$ and $G=200$, we simulated $\mathbf{X}$ and $\bm{\beta}$ as in Experiment 1 of Section 6.2. Then we simulated binary responses $y_{i}\mid\mathbf{x}_{i}\sim\text{Bernoulli}(1/(1+\exp(-\theta_{i}))),i=1,\ldots,n$, where $\log(\theta_{i}/(1-\theta_{i}))=\mathbf{x}_{i}^{\top}\bm{\beta}$.
  • Experiment 10 (grouped Poisson regression). With $n=300$ and $G=200$, we simulated $\mathbf{X}$ and $\bm{\beta}$ as in Experiment 5 of Section 6.2. Then we simulated count responses $y_{i}\mid\mathbf{x}_{i}\sim\text{Poisson}(\exp(\theta_{i})),i=1,\ldots,n$, where $\log(\theta_{i})=\mathbf{x}_{i}^{\top}\bm{\beta}$.
It should be noted that in Experiments 9 and 10, the group sizes ranged from three to five. Thus, Experiments 9 and 10 dealt with the $p>n$ scenario where $p$ could range from 600 to 1000.
After generating the data, we used the Gibbs sampling algorithm of Section 5.3 to draw posterior samples of the regression coefficients $\bm{\beta}=(\beta_{1},\ldots,\beta_{p})^{\top}$. We used the same hyperparameters as those suggested in Section 5.2, except we fixed $\lambda_{0}=5$. We ran the Gibbs sampler for a total of 3000 iterations, discarding the first 1000 samples as burnin. The remaining 2000 samples were used to approximate the marginal posterior distributions for the regression coefficients. To assess the quality of the SSGL posterior approximation, we recorded the following metrics:
@@
    the coverage probability (CP) of the 95% posterior credible intervals (CIs), i.e. the proportion of CIs for the regression coefficients $\beta_{j},j=1,\ldots,p$, that contained the true $\beta_{j}$. These intervals were constructed using the 0.025 and 0.975 sample quantiles $[q_{0.025}(\beta_{j}),q_{0.975}(\beta_{j})]$ of the saved 2000 MCMC samples for each $\beta_{j}$.
@@
    the average width of the 95% CIs for all regression coefficients and the average width of the 95% CIs for the nonzero regression coefficients, denoted as Width and $\text{Width}_{\text{nonzero}}$ respectively;
@@
We repeated Experiments 9 and 10 for 200 replications. Table 3 reports the CP, Width, $\text{Width}_{\text{nonzero}}$, ESS, $\text{MCSE}(q_{0.025})$ and $\text{MCSE}(q_{0.975})$ averaged across these 200 replicates. We see that for both logistic and Poisson regression, the 95% CIs largely covered the true regression coefficients in $\bm{\beta}$. Meanwhile, the Width and $\text{Width}_{\text{nonzero}}$ reported in Table 3 show that the 95% CIs were not overwhelmingly conservative. Thus, the high CP was not a consequence of the CIs being too wide. Our results verify the usefulness of our SSGL Gibbs sampling algorithm for quantifying uncertainty of $\bm{\beta}$.
@@
CP 0.997 (0.004) 0.993 (0.005)
Width 0.871 (0.015) 1.51 (0.646)
$\text{Width}_{\text{nonzero}}$ 2.98 (0.375) 2.74 (1.22)
ESS 1619.78 (18.94) 672.86 (87.68)
$\text{MCSE}(q_{0.025})$ 0.014 (0.002) 0.051 (0.049)
$\text{MCSE}(q_{0.975})$ 0.014 (0.002) 0.033 (0.032)
Table 3 shows that in Experiment 9 (grouped logistic regression), the average ESS for the 2000 saved MCMC samples 1619.78, suggesting high efficiency of our Gibbs sampling algorithm. The $\text{MCSE}(q_{0.025})$ and $\text{MCSE}(q_{0.975})$ in Experiment 9 also suggested fairly high precision of the MCMC estimates for the 95% CI endpoints. In Experiment 10 (grouped Poisson regression), the average ESS for the 2000 saved MCMC samples was 672.86, and the average $\text{MCSE}(q_{0.025})$ and $\text{MCSE}(q_{0.975})$ were higher than those in Experiment 9. This suggests that there was some loss of efficiency from using the approximation (5.11) for the Poisson log-likelihood in our Gibbs sampler. However, the average ESS and the precisions of our MCMC estimates were still acceptable in Experiment 10.
Figure 1 plots the marginal posterior distributions for the regression coefficients from one of the nonzero groups in one replication of Experiment 9 (top panel) and one replication of Experiment 10 (bottom panel). The true parameter values are plotted as red vertical lines. Figure 1 shows that in both grouped logistic regression and grouped Poisson regression, the SSGL posterior was able to capture the ground truth.
@@
Figure 1: Estimated marginal posterior densities for the regression coefficients from one of the nonzero group in one replication of Experiment 9 (top panel) and one replication of Experiment 10 (bottom panel). The true parameter values are plotted as red vertical lines.
@@ 7 Application to HIV drug resistance data
8 Discussion
In this paper, we have extended the SSGL prior of Bai et al. [3] from Gaussian linear regression models to GLMs (1.3) under the general exponential family (1.1). This enables the SSGL (2.2) to be applied to binary regression, Poisson regression, negative binomial regression, and gamma regression, among others.
In contrast to high-dimensional Bayesian linear regression models, theoretical developments for high-dimensional Bayesian GLMs have been much slower. Our work builds upon and goes beyond those of Jiang [21], Jeong and Ghosal [20], and Tang and Martin [37] in several ways. First, we conducted our analysis under a continuous heavy-tailed spike-and-slab prior (2.2), rather than a point mass spike-and-slab prior. Second, we considered the more general grouped GLM setting (1.3). By treating unstructured GLMs (1.2) as a special case of (1.3), we also obtained the first theoretical results for the SSL prior (2.3) of Ročková and George [35] in high-dimensional GLMs. Finally, we analyzed both the MAP estimator (2.6) and the full posterior distribution (2.4), confirming that the rate of convergence is the same for both. This is in stark contrast to the original group lasso (2.7). We showed via extensive simulation studies and an analysis of an HIV drug resistance dataset that the SSGL performs very well in grouped GLMs when $G<n$ and $G>n$.
Although we have addressed Bayesian estimation and variable selection in high-dimensional GLMs, there are still some open theoretical questions. Theorem 3 implies that the full SSGL posterior provides valid inference in terms of the size of credible sets. Apart from the size of the posterior credible sets, however, theoretical questions about the asymptotic coverage of Bayesian credible sets and/or a Bernstein-von Mises theorem for the limiting shape of the posterior are also important questions that have not been addressed in the literature. This is an important future research direction for high-dimensional Bayesian GLMs.
@@ Acknowledgments
The author is grateful to two anonymous reviewers whose thoughtful feedback helped to greatly improve this paper. The author also thanks Dr. Seonghyun Jeong for helpful discussions which led to derivations of the theoretical results in this paper.
@@ Appendix A Proofs for Section 3
Under the objective function (2.5), it can then be seen from (1.5) and (A.3) that any local maximizer of (2.5) must satisfy
$\mathbf{X}_{g}^{\top}\textrm{diag}\{\xi^{\prime}(\mathbf{X}\bm{\beta})\}\left(\mathbf{Y}-{\color[rgb]{0,0,0}b^{\prime}(\xi(\mathbf{X}\bm{\beta}))}\right)-\lambda_{\theta}^{\star}(\bm{\beta}_{g})\partial\lVert\bm{\beta}_{g}\rVert_{2}=\bm{0}_{m_{g}}\textrm{for all }g\in\{1,\ldots,G\},$ (A.5)
@@
$Q_{1}(\bm{\beta}_{0}+\epsilon_{n}\bm{\Delta})-Q_{1}(\bm{\beta}_{0})$
$\qquad=\left[\ell_{n}(\bm{\beta}_{0}+\epsilon_{n}\bm{\Delta})-\ell_{n}(\bm{\beta}_{0})\right]+\sum_{g=1}^{s_{0}}\left[\textrm{pen}(\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g})-\textrm{pen}(\bm{\beta}_{0g})\right]$
=$\qquad\overset{\Delta}{=}(\textrm{I})+(\textrm{II})<0,$ (A.10)
@@
$\qquad\qquad\leq\frac{d_{2}^{2}n}{A^{2}s_{0}\log G}\mathbb{E}\left\{\sum_{g=1}^{s_{0}}\sum_{k=1}^{m_{g}}\left[\frac{1}{n}\sum_{i=1}^{n}x_{igk}\left(y_{i}-{\color[rgb]{0,0,0}b^{\prime}(\xi(\mathbf{x}_{i}^{\top}\bm{\beta}_{0}))}\right)\right]^{2}\right\}$
$\qquad\qquad=\frac{d_{2}^{2}}{A^{2}s_{0}\log G}\textrm{tr}\left(\frac{1}{n}\mathbf{X}_{S_{0}}^{\top}{\color[rgb]{0,0,0}\bm{\Omega}(\bm{\beta}_{0})}\mathbf{X}_{S_{0}}\right)$
$\qquad\qquad\leq\frac{d_{2}^{2}}{A^{2}s_{0}\log G}\left(\sum_{g=1}^{s_{0}}m_{g}\right)\lambda_{\max}\left(\frac{1}{n}\mathbf{X}_{S_{0}}^{\top}{\color[rgb]{0,0,0}\bm{\Omega}(\bm{\beta}_{0})}\mathbf{X}_{S_{0}}\right)$
$\qquad\qquad\leq\frac{d_{2}^{2}s_{0}m_{\max}}{A^{2}s_{0}\log G}\lambda_{\max}\left(\frac{1}{n}\mathbf{X}_{S_{0}}^{\top}{\color[rgb]{0,0,0}\bm{\Omega}(\bm{\beta}_{0})}\mathbf{X}_{S_{0}}\right)$
$\qquad\qquad{\color[rgb]{0,0,0}\lesssim}\frac{d_{2}^{2}\log n}{A^{2}\log G}\lambda_{\max}\left(\frac{1}{n}\mathbf{X}_{S_{0}}^{\top}{\color[rgb]{0,0,0}\bm{\Omega}(\bm{\beta}_{0})}\mathbf{X}_{S_{0}}\right)\rightarrow 0~{}~{}{\color[rgb]{0,0,0}\textrm{as}~{}~{}n\rightarrow\infty},$ (A.12)
@@
$\textrm{pen}(\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g})-\textrm{pen}(\bm{\beta}_{0g})$
$\qquad=\lambda_{1}\left\{\lVert\bm{\beta}_{0g}\rVert_{2}-\lVert\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g}\rVert_{2}\right\}+\left[\log\left(p_{\theta}^{\star}(\bm{\beta}_{0g})\right)-\log\left(p_{\theta}^{\star}(\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g})\right)\right]$
$\qquad\leq\lambda_{1}\epsilon_{n}\lVert\bm{\Delta}_{g}\rVert_{2}-\left[\log(p_{\theta}^{\star}(\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g}))-\log(p_{\theta}^{\star}(\bm{\beta}_{0g}))\right]$
$\qquad\leq\lambda_{1}\epsilon_{n}\lVert\bm{\Delta}_{g}\rVert_{2}-\epsilon_{n}\left(\frac{\partial p_{\theta}^{\star}(\widetilde{\bm{\beta}}_{g})}{\partial\widetilde{\bm{\beta}}_{g}}\right)^{\top}\bm{\Delta}_{g},$
@@
$\qquad\leq\epsilon_{n}\lVert\bm{\Delta}_{g}\rVert_{2}\left\{\lambda_{1}+\bigg{\lVert}\frac{\partial p_{\theta}^{\star}(\widetilde{\bm{\beta}}_{g})}{\partial\widetilde{\bm{\beta}}_{g}}\bigg{\rVert}_{2}\right\},$ (A.14)
@@
$p_{\theta}^{\star}(\bm{\beta}_{m_{g}})=\frac{1}{1+\left(\frac{1-\theta}{\theta}\right)\left(\frac{\lambda_{0}}{\lambda_{1}}\right)^{m_{g}}\exp\left[-(\lambda_{0}-\lambda_{1})\lVert\bm{\beta}_{g}\rVert_{2}\right]}:=\frac{1}{1+c(\bm{\beta}_{g})},$ (A.15)
@@
$(\textrm{II})~{}{\color[rgb]{0,0,0}\lesssim}~{}\epsilon_{n}\sum_{g=1}^{s_{0}}(1+\lambda_{1})\lVert\bm{\Delta}_{g}\rVert_{2}$ $\leq s_{0}\epsilon_{n}(1+\lambda_{1})\max_{1\leq g\leq s_{0}}\lVert\bm{\Delta}_{g}\rVert_{2}$
$\leq s_{0}\epsilon_{n}(1+\lambda_{1})\lVert\bm{\Delta}_{S_{0}}\rVert_{2}.$ (A.16)
Combining the upper bounds in (A.13) and (A), it is clear that
@@
$\lambda_{\max}\left\{\nabla^{2}\ell_{n}(\widehat{\bm{\beta}}_{g})+\nabla^{2}\textrm{pen}(\widehat{\bm{\beta}}_{g})\right\}<0~{}~{}\textrm{for all }g\in S_{0},$ (A.17)
where $\textrm{pen}(\bm{\beta}_{g})$ is as in (A.1), and
$\lVert\mathbf{X}_{g}^{\top}\textrm{diag}(\xi^{\prime}(\mathbf{X}\widehat{\bm{\beta}}))(\mathbf{Y}-{\color[rgb]{0,0,0}b^{\prime}(\xi(\mathbf{X}\bm{\beta})))}\rVert_{2}<\lambda_{\bm{\theta}}^{\star}(\bm{0}_{m_{g}})~{}~{}\textrm{for all}g\in S_{0}^{c}.$ (A.18)
@@
We first prove (A.17). We have that
$\nabla^{2}\ell_{n}(\widehat{\bm{\beta}}_{g})+\nabla^{2}\textrm{pen}(\widehat{\bm{\beta}}_{g})$
$\qquad=-\mathbf{X}_{g}^{\top}\bm{\Sigma}(\widehat{\bm{\beta}})\mathbf{X}_{g}+p_{\theta}^{\star}(\widehat{\bm{\beta}}_{g})[1-p_{\theta}^{\star}(\widehat{\bm{\beta}}_{g})](\lambda_{0}-\lambda_{1})^{2}\left[\frac{1}{\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}}\mathbf{I}_{m_{g}}-\frac{\widehat{\bm{\beta}}_{g}\widehat{\bm{\beta}}_{g}^{\top}}{\lVert\bm{\beta}_{g}\rVert_{2}^{3}}\right].$
Thus,
$\lambda_{\max}\left\{\nabla^{2}\ell_{n}(\widehat{\bm{\beta}}_{g})+\nabla^{2}\textrm{pen}(\widehat{\bm{\beta}}_{g})\right\}$
$\leq-\lambda_{\min}\left(-\mathbf{X}_{g}^{\top}\bm{\Sigma}(\widehat{\bm{\beta}})\mathbf{X}_{g}\right)+\frac{(\lambda_{0}-\lambda_{1})^{2}}{4}\left[\lambda_{\max}\left(\frac{1}{\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}}\mathbf{I}_{m_{g}}\right)-\lambda_{\min}\left(\frac{\widehat{\bm{\beta}}_{g}\widehat{\bm{\beta}}_{g}^{\top}}{\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{3}}\right)\right]$
${\color[rgb]{0,0,0}\lesssim}-n+\frac{(\lambda_{0}-\lambda_{1})^{2}}{4}\left[\frac{1}{\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}}-\frac{1}{\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}}\right]=-n<0,$
where we used Weyl’s inequality, Assumption (A3)(ii), the fact that $\lambda_{\max}(-\mathbf{A})=-\lambda_{\min}(\mathbf{A})$ for a square matrix $\mathbf{A}$, and the fact that $\widehat{\bm{\beta}}_{g}\widehat{\bm{\beta}}_{g}^{\top}$ has the same eigenvalue as $\widehat{\bm{\beta}}_{g}^{\top}\widehat{\bm{\beta}}_{g}=\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{2}$. Thus, we have shown that (A.17) holds for large $n$.

### v8

@@ 6.4 Fully Bayesian inference
  • Experiment 9 (grouped logistic regression). With $n=400$ and $G=250$ groups of size four (i.e. $p=1000$), we simulated $\mathbf{X}$ as in Experiment 1 of Section 6.2. We randomly chose $s_{0}=10$ of the vectors $\bm{\beta}_{g}$ to be nonzero with $\bm{\beta}_{g}=(-1,-0.8,0.8,1)^{\top}$, and the remaining groups were set to zero. Then we simulated binary responses $y_{i}\mid\mathbf{x}_{i}\sim\text{Bernoulli}(1/(1+\exp(-\theta_{i}))),i=1,\ldots,n$, where $\log(\theta_{i}/(1-\theta_{i}))=\mathbf{x}_{i}^{\top}\bm{\beta}$.
  • Experiment 10 (grouped Poisson regression). With $n=300$ and $G=200$ groups of size four (i.e. $p=800$), we simulated $\mathbf{X}$ as in Experiment 5 of Section 6.2. We randomly chose $s_{0}=10$ of the vectors $\bm{\beta}_{g}$ to be nonzero with $\bm{\beta}_{g}=(-0.9,-0.7,0.7,0.9)^{\top}$, and the remaining groups were set to zero. Then we simulated count responses $y_{i}\mid\mathbf{x}_{i}\sim\text{Poisson}(\exp(\theta_{i})),i=1,\ldots,n$, where $\log(\theta_{i})=\mathbf{x}_{i}^{\top}\bm{\beta}$.
After generating the data, we used the Gibbs sampling algorithm of Section 5.3 to draw posterior samples. We used the same hyperparameters as those suggested in Section 5.2. In particular, the spike parameter $\lambda_{0}$ was first tuned for the SSGL MAP estimator using the EM algorithm and then fixed at this value for the MCMC algorithm. We ran the Gibbs sampler for a total of 3000 iterations, discarding the first 1000 samples as burnin. The remaining 2000 samples were used to approximate the marginal posterior distributions for the regression coefficients. To assess the quality of the SSGL posterior approximation, we recorded the following metrics:
@@
    the coverage probability (CP) of the 95% posterior credible intervals (CIs) for the 40 true non-null coefficients in $\bm{\beta}$, i.e. the proportion of CIs for the nonzero regression coefficients $\{\beta_{j}:\beta_{j}\neq 0\}$ that contained the true $\beta_{j}$. These intervals were constructed using the 0.025 and 0.975 sample quantiles $[q_{0.025}(\beta_{j}),q_{0.975}(\beta_{j})]$ of the saved 2000 MCMC samples for each $\beta_{j}$.
@@
    the average width of the 95% CIs for the 40 non-null regression coefficients;
@@
We repeated Experiments 9 and 10 for 200 replications. Table 3 reports the CP, Width, ESS, $\text{MCSE}(q_{0.025})$ and $\text{MCSE}(q_{0.975})$ averaged across these 200 replicates. We see that in both experiments, the 95% CIs covered the true non-null regression coefficients in $\bm{\beta}$ at close to the nominal rate. Table 3 also shows that the average widths of the 95% CIs for the true nonzero coefficients were not overwhelmingly conservative. Thus, this coverage was not achieved by the CIs being too wide. Our results verify the usefulness of our SSGL Gibbs sampling algorithm for quantifying uncertainty of $\bm{\beta}$.
@@
CP 0.960 (0.032) 0.937 (0.039)
Width 2.44 (0.238) 1.81 (0.048)
ESS 1523.29 (13.91) 1255.78 (53.92)
$\text{MCSE}(q_{0.025})$ 0.034 (0.007) 0.033 (0.004)
$\text{MCSE}(q_{0.975})$ 0.030 (0.008) 0.033 (0.008)
Table 3 shows that in Experiment 9 (grouped logistic regression), the average ESS for the 2000 saved MCMC samples was 1523.29, suggesting high efficiency of our Gibbs sampling algorithm. In Experiment 10 (grouped Poisson regression), the average ESS for the 2000 saved MCMC samples was slightly lower (1255.78), and the CIs had average coverage slightly below the nominal level. This suggests that there was some loss of efficiency from using the approximation (5.11) for the Poisson log-likelihood in our Gibbs sampler. However, the CP and ESS were still acceptable in Experiment 10. Finally, in both experiments, $\text{MCSE}(q_{0.025})$ and $\text{MCSE}(q_{0.975})$ indicated an acceptable level of precision for the 0.025 and 0.975 MCMC sample quantiles.
Figure 1: Estimated marginal posterior densities for four regression coefficients from one replication of Experiment 9 (top panel) and four regression coefficients from one replication of Experiment 10 (bottom panel). The true parameter values are plotted as solid red vertical lines, and the SSGL MAP estimators are plotted as dashed blue lines. The left four plots depict the results for four nonzero coefficients, while the right four plots depict the results for four null coefficients where the SSGL MAP estimate is also exactly zero.
Figure 1 plots the kernel-smoothed marginal posterior densities for four regression coefficients from one replication of Experiment 9 (top panel) and four regression coefficients from one replication of Experiment 10 (bottom panel). The true parameter values are plotted as solid red vertical lines, and the SSGL MAP estimates are plotted as dashed vertical lines. Figure 1 shows that in both grouped logistic regression and grouped Poisson regression, the SSGL posterior was able to capture the ground truth. Moreover, the SSGL MAP estimates were close to their true values for the true nonzero coefficients, and they were exactly equal to zero for the true null coefficients. This is due to the exact sparsity of the SSGL MAP estimate (2.6).
@@ Acknowledgments
The author is grateful to the anonymous reviewers whose thoughtful feedback helped to greatly improve this paper. The author also thanks Dr. Seonghyun Jeong for helpful discussions which led to derivations of the theoretical results in this paper.
@@ Appendix A Proofs for Section 3
Under the objective function (2.5), it can then be seen from (1.5) and (A.3) that any local maximizer $\widehat{\bm{\beta}}$ of (2.5) must satisfy
$\mathbf{X}_{g}^{\top}\textrm{diag}\{\xi^{\prime}(\mathbf{X}{\color[rgb]{0,0,0}\widehat{\bm{\beta}}})\}\left(\mathbf{Y}-{\color[rgb]{0,0,0}b^{\prime}(\xi(\mathbf{X}{\color[rgb]{0,0,0}\widehat{\bm{\beta}}}))}\right)-\lambda_{\theta}^{\star}({\color[rgb]{0,0,0}\widehat{\bm{\beta}}_{g}})\partial\lVert\bm{\beta}_{g}\rVert_{2}=\bm{0}_{m_{g}}\textrm{for all }g\in\{1,\ldots,G\},$ (A.5)
@@
$Q_{1}(\bm{\beta}_{0}+\epsilon_{n}\bm{\Delta})-Q_{1}(\bm{\beta}_{0})$ $=\left[\ell_{n}(\bm{\beta}_{0}+\epsilon_{n}\bm{\Delta})-\ell_{n}(\bm{\beta}_{0})\right]+\sum_{g=1}^{s_{0}}\left[\textrm{pen}(\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g})-\textrm{pen}(\bm{\beta}_{0g})\right]$
=$\overset{\Delta}{=}(\textrm{I})+(\textrm{II})<0,$ (A.10)
@@
$\leq\frac{d_{2}^{2}n}{A^{2}s_{0}\log G}\mathbb{E}\left\{\sum_{g=1}^{s_{0}}\sum_{k=1}^{m_{g}}\left[\frac{1}{n}\sum_{i=1}^{n}x_{igk}\left(y_{i}-{\color[rgb]{0,0,0}b^{\prime}(\xi(\mathbf{x}_{i}^{\top}\bm{\beta}_{0}))}\right)\right]^{2}\right\}$
$=\frac{d_{2}^{2}}{A^{2}s_{0}\log G}\textrm{tr}\left(\frac{1}{n}\mathbf{X}_{S_{0}}^{\top}{\color[rgb]{0,0,0}\bm{\Omega}(\bm{\beta}_{0})}\mathbf{X}_{S_{0}}\right)$
$\leq\frac{d_{2}^{2}}{A^{2}s_{0}\log G}\left(\sum_{g=1}^{s_{0}}m_{g}\right)\lambda_{\max}\left(\frac{1}{n}\mathbf{X}_{S_{0}}^{\top}{\color[rgb]{0,0,0}\bm{\Omega}(\bm{\beta}_{0})}\mathbf{X}_{S_{0}}\right)$
$\leq\frac{d_{2}^{2}s_{0}m_{\max}}{A^{2}s_{0}\log G}\lambda_{\max}\left(\frac{1}{n}\mathbf{X}_{S_{0}}^{\top}{\color[rgb]{0,0,0}\bm{\Omega}(\bm{\beta}_{0})}\mathbf{X}_{S_{0}}\right)$
${\color[rgb]{0,0,0}\lesssim}\frac{d_{2}^{2}\log n}{A^{2}\log G}\lambda_{\max}\left(\frac{1}{n}\mathbf{X}_{S_{0}}^{\top}{\color[rgb]{0,0,0}\bm{\Omega}(\bm{\beta}_{0})}\mathbf{X}_{S_{0}}\right)\rightarrow 0~{}~{}{\color[rgb]{0,0,0}\textrm{as}~{}~{}n\rightarrow\infty},$ (A.12)
@@
$\textrm{pen}(\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g})-\textrm{pen}(\bm{\beta}_{0g})$ $=\lambda_{1}\left\{\lVert\bm{\beta}_{0g}\rVert_{2}-\lVert\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g}\rVert_{2}\right\}+\left[\log\left(p_{\theta}^{\star}(\bm{\beta}_{0g})\right)-\log\left(p_{\theta}^{\star}(\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g})\right)\right]$
$\leq\lambda_{1}\epsilon_{n}\lVert\bm{\Delta}_{g}\rVert_{2}-\left[\log(p_{\theta}^{\star}(\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g}))-\log(p_{\theta}^{\star}(\bm{\beta}_{0g}))\right]$
$\leq\lambda_{1}\epsilon_{n}\lVert\bm{\Delta}_{g}\rVert_{2}-\epsilon_{n}\left(\frac{\partial p_{\theta}^{\star}(\widetilde{\bm{\beta}}_{g})}{\partial\widetilde{\bm{\beta}}_{g}}\right)^{\top}\bm{\Delta}_{g},$
@@
$\leq\epsilon_{n}\lVert\bm{\Delta}_{g}\rVert_{2}\left\{\lambda_{1}+\bigg{\lVert}\frac{\partial p_{\theta}^{\star}(\widetilde{\bm{\beta}}_{g})}{\partial\widetilde{\bm{\beta}}_{g}}\bigg{\rVert}_{2}\right\},$ (A.14)
@@
$p_{\theta}^{\star}({\color[rgb]{0,0,0}\bm{\beta}_{g}})=\frac{1}{1+\left(\frac{1-\theta}{\theta}\right)\left(\frac{\lambda_{0}}{\lambda_{1}}\right)^{m_{g}}\exp\left[-(\lambda_{0}-\lambda_{1})\lVert\bm{\beta}_{g}\rVert_{2}\right]}:=\frac{1}{1+c(\bm{\beta}_{g})},$ (A.15)
@@
$(\textrm{II})~{}{\color[rgb]{0,0,0}\lesssim}~{}\epsilon_{n}\sum_{g=1}^{s_{0}}(1+\lambda_{1})\lVert\bm{\Delta}_{g}\rVert_{2}$ $\leq s_{0}\epsilon_{n}(1+\lambda_{1})\max_{1\leq g\leq s_{0}}\lVert\bm{\Delta}_{g}\rVert_{2}\leq s_{0}\epsilon_{n}(1+\lambda_{1})\lVert\bm{\Delta}_{S_{0}}\rVert_{2}.$ (A.16)
Combining the upper bounds in (A.13) and (A.16), it is clear that
@@
$\lambda_{\max}\left\{{\color[rgb]{0,0,0}\nabla_{g}^{2}\ell_{n}(\widehat{\bm{\beta}})}+\nabla^{2}\textrm{pen}(\widehat{\bm{\beta}}_{g})\right\}<0~{}~{}\textrm{for all }{\color[rgb]{0,0,0}\widehat{\bm{\beta}}_{g}\neq\bm{0}_{m_{g}}},g\in S_{0},$ (A.17)
where $\nabla_{g}^{2}$ denotes the second partial derivative with respect to $\widehat{\bm{\beta}}_{g}$, $\textrm{pen}(\bm{\beta}_{g})$ is the penalty function in (A.1), and
$\lVert\mathbf{X}_{g}^{\top}\textrm{diag}(\xi^{\prime}(\mathbf{X}\widehat{\bm{\beta}}))[\mathbf{Y}-{\color[rgb]{0,0,0}b^{\prime}(\xi({\color[rgb]{0,0,0}\mathbf{X}\widehat{\bm{\beta}}}))]}\rVert_{2}<\lambda_{\bm{\theta}}^{\star}(\bm{0}_{m_{g}})~{}~{}\textrm{for all }g\in S_{0}^{c}.$ (A.18)
@@
We first prove (A.17). For $\widehat{\bm{\beta}}_{g}\neq\bm{0}_{m_{g}}$, we have by (1.6) and the chain rule for second derivatives (with $f(\mathbf{u})=\text{pen}(\mathbf{u})$ and $\mathbf{u}=g(\mathbf{v})=\lVert\mathbf{v}\rVert_{2}$) that
${\color[rgb]{0,0,0}\nabla_{g}^{2}}\ell_{n}(\widehat{\bm{\beta}})+\nabla^{2}\textrm{pen}(\widehat{\bm{\beta}}_{g})$
$\qquad=-\mathbf{X}_{g}^{\top}\bm{\Sigma}(\widehat{\bm{\beta}})\mathbf{X}_{g}{\color[rgb]{0,0,0}~{}+~{}p_{\theta}^{\star}(\widehat{\bm{\beta}}_{g})[1-p_{\theta}^{\star}(\widehat{\bm{\beta}}_{g})](\lambda_{0}-\lambda_{1})^{2}\frac{\widehat{\bm{\beta}}_{g}\widehat{\bm{\beta}}_{g}^{\top}}{\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{2}}}{\color[rgb]{0,0,0}-~{}\lambda_{\theta}^{\star}(\widehat{\bm{\beta}}_{g})\left[\frac{1}{\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}}\mathbf{I}_{m_{g}}-\frac{\widehat{\bm{\beta}}_{g}\widehat{\bm{\beta}}_{g}^{\top}}{\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{3}}\right].}$ (A.19)
Note that $\lambda_{\max}(\lVert\bm{\widehat{\beta}}_{g}\rVert_{2}^{-2}\bm{\widehat{\beta}}_{g}\bm{\widehat{\beta}}_{g}^{\top})=1$, since $\bm{\widehat{\beta}}_{g}\bm{\widehat{\beta}}_{g}^{\top}$ has the same nonzero eigenvalue as $\bm{\widehat{\beta}}_{g}^{\top}\bm{\widehat{\beta}}_{g}=\lVert\bm{\widehat{\beta}}_{g}\rVert_{2}^{2}$. Meanwhile, the only eigenvalue of $\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{-1}\mathbf{I}_{m_{g}}$ is $\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{-1}$, and the only eigenvalues of $\lVert\bm{\widehat{\beta}}_{g}\rVert_{2}^{-3}\bm{\widehat{\beta}}_{g}\bm{\widehat{\beta}}_{g}^{\top}$ are 0 and $\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{-1}$. Consequently, $-\lambda_{\min}(\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{-1}\mathbf{I}_{m_{g}}-\lVert\bm{\widehat{\beta}}_{g}\rVert_{2}^{-3}\bm{\widehat{\beta}}_{g}\bm{\widehat{\beta}}_{g}^{\top})\leq 0$. Using Weyl’s inequality and the fact that $\lambda_{\max}(-\mathbf{A})=-\lambda_{\min}(\mathbf{A})$ for a square matrix $\mathbf{A}$, it follows from (A) that
$\lambda_{\max}\left\{{\color[rgb]{0,0,0}\nabla_{g}^{2}}\ell_{n}(\widehat{\bm{\beta}})+\nabla^{2}\textrm{pen}(\widehat{\bm{\beta}}_{g})\right\}$ $\leq-\lambda_{\min}\left(\mathbf{X}_{g}^{\top}\bm{\Sigma}(\widehat{\bm{\beta}})\mathbf{X}_{g}\right)+~{}p_{\theta}^{\star}(\widehat{\bm{\beta}}_{g})[1-p_{\theta}^{\star}(\widehat{\bm{\beta}}_{g})](\lambda_{0}-\lambda_{1})^{2}$
[…]
$\leq-n+c(\widehat{\bm{\beta}}_{g})(\lambda_{0}-\lambda_{1})^{2}$
$\lesssim-n+1<0,$
for sufficiently large $n$, where we used Assumption (A3)(ii) in the third line and the fact that given our choices of hyperparameters for $(\theta,\lambda_{0},\lambda_{1})$ in Theorem 1, $c(\bm{\beta}_{g})(\lambda_{0}-\lambda_{1})^{2}\prec 1$ for any $\bm{\beta}_{g}\neq\bm{0}_{m_{g}}$ in the last line. Thus, we have shown that (A.17) holds for large $n$.